OpenStudy (anonymous):

Please hELP WILL bECOME FAN!!!! AND gIVE mEDAL iF rIGHT!!! 27. 1 and 2 are a linear pair. m1= x – 35, and m 2 = x + 83. Find the measure of each angle. (1 point) 1 = 31°, 2 = 149° 1 = 66°, 2 = 114° 1 = 31°, 2 = 159° 1 = 66°, 2 = 124°

OpenStudy (anonymous):



OpenStudy (anonymous):

oH SO 118 - 180 = 62 for both X's ?

OpenStudy (anonymous):

Wait.. when you solved the equation, what did you get for X?

OpenStudy (anonymous):

62

OpenStudy (anonymous):

..How did you get that?

OpenStudy (anonymous):

I added 35 +83 = 118 - 180 = 62 / 2

OpenStudy (anonymous):

No heres how you do it: x - 35 + x+ 83 = 180 (combine like terms) 2x + 48 = 180 -48 -48 -------------- 2x = 132 --------- 2 x = 66

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Now use that x value, and plug it into the problem for the angle 1. That will be the answer for the m of angle 1. Do the same for angle 2.

OpenStudy (anonymous):

oh so 31 and 149

OpenStudy (anonymous):

Yep, thats perfect.

OpenStudy (anonymous):

Could you help me with a few more questions? ANd thank you so much

OpenStudy (anonymous):

Sure, close this post and open another one.